OpenStudy (anonymous):

If the common difference in an arithmetic sequence is –3, and the first term is 7, what is the 14th term?

OpenStudy (imstuck):

Do you know the equation for finding the nth term in an arithmetic series?

OpenStudy (anonymous):

Between the 1st term and the 14th term there are (14 -1 = 13) common differences. The nth term is the first term + (n-1)* common difference Therefore the 14th term is 3 + (13*3) = 3 + 39 = 42 check 3,6,9,12,15,18,21,24,27,30,33,36,39,42 try to follow this equation .